10:00: cal unit deployed

We start testing the new RECs: we load GLAO2 PYR BIN2 generated by TM

TN Notes
20150701_182848 GLAO2, pyr bin2
20150701_182855 Mirror PANIC

We go to open the Adsec cabinet door: telescope goes to horizon.

TN Notes
20150701_195443 GLAO2 PYR BIN2
20150701_195503 "
20150701_195720 GLAO1 PYR BIN2
20150701_195738 "
20150701_195914 TCoG48 PYR BIN2
20150701_195932 "

We move to QC hot spot detection on CCD47. We close the TT loop on QC, FW1 with dichroic 700-1000, FW2 QC mirror. Then we pause the loop and we exchange the filter on FW1 moving to 50-50:

Board X Y Z
  -70.3 -47.4 39.3

CCD47 BIN4 hot pixel: [127; 133] with 50-50 on FW1

CCD47 BIN4 hot pixel: [129; 127] with 700-1000 on FW1

We try to close the loop on LUCI N30 to check the LGSW performance with pyramid copointing. Pyr piezo is at [-2.20; 0.06].

-- MarcoBonaglia - 05 Aug 2015
Topic revision: r1 - 05 Aug 2015, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback